# A detailed study on ergodic to non-ergodic phase transition in the dissipative anisotropic Dicke model

### v1

@@ Abstract
Ergodic-non-ergodic phase transition is one of the paramount features of the anisotropic Dicke model. Here, we have thoroughly examined the effect of the dissipation by analyzing both the eigenvalue and eigenvector properties of the Liouvillian with the aid of the scaling of Liouvilian gap, and the average participation ratio. We show that the properties of the eigenvectors of Liouvillian are consistent with those of the eigenvalues, revealing a phase diagram, which has similarities to the non-ergodic to ergodic transition in the closed undriven system. We also uncover that the Liouvillian gap is independent of system size in the non-ergodic phase whereas in the ergodic phase, it scales with the atom number as: $N^{-z}$ where $10<z<$ 1. Moreover, we extend our analysis to the driven case where a Thue-Morse quasi-periodic sequence is applied and observe that the boson dissipation plays a pivotal role in stabilizing the prethermal plateau. Our investigation indicates that a non-ergodic phase is more favorable than the ergodic phase in the presence of bosonic dissipation.
@@ I Introduction
Exploring the many-body quantum dynamics through atom-photon interaction is the heart of the atomic, condensed matter, and ultra-cold quantum physics Cohen-Tannoudji et al. (1998); Gutzler et al. (2021); Chang et al. (2018). Experiments using cavity QED setup provide an ideal platform to investigate the underlying microscopic properties e.g., the collective effects, various phase transitions, Bose-Einstein condensation, etc Walther et al. (2006); Reiserer and Rempe (2015); Vaidya et al. (2018). From a theoretical perspective, the Dicke model (DM) is an iconic model for understanding the above aspects of such systems Garraway (2011); Kirton et al. (2019); Villaseñor et al. (2024).
@@
The Dicke model describes the interaction between $N$ ensemble of atoms with a single mode bosonic field via a dipolar coupling strength  Dicke (1954); Lambert et al. (2004); Emary and Brandes (2003a, b). In the thermodynamic limit ($N\to\infty$) the model shows a quantum phase transition from normal to superradiant phase Dicke (1954); Chávez-Carlos et al. (2016); Kirton and Keeling (2018); Kirton et al. (2019); Das and Sharma (2022). Along with it, such a model also shows the excited state phase transition(ESQPT), and thermal phase transition(TPT) under different circumstances. A generalized version of the Dicke model namely the anisotropic Dicke model (ADM) Buijsman et al. (2017); Kloc et al. (2017); Shapiro et al. (2020); Das et al. (2023a), where the coupling strengths correspond to the rotating and counter-rotating terms are considered to be different. Due to the asymmetry in the coupling strengths, ADM shows novel features in addition to the well-known existing properties, i.e., ergodic to non-ergodic transition (ENET) Buijsman et al. (2017); Hu and Wan (2021). There exists a one-to-one correspondence of the non-ergodic phase to the integrability of the system. Hence, a detailed study on the non-ergodic phase will be useful in understanding the following emerging phenomena, e.g., many-body localization, quantum scars, prethermalization, etc Polkovnikov et al. (2011); Mori et al. (2018); Serbyn et al. (2021). On the other hand, the ergodic phase connects to the non-integrable phase, a path to study the quantum chaos D’Alessio et al. (2016); Borgonovi et al. (2016). Generally, the question arises regarding the behavior of the above-mentioned phases in the presence of dissipation.
In most cases, a leaky cavity is considered to introduce the dissipative effects on the atom-photon interaction Bhaseen et al. (2012); Kirton et al. (2019); Kirton and Keeling (2017, 2018). It was shown that such an effect is important for understanding the self-organization and discrete time-crystalline order in this model Keeling et al. (2010); Gong et al. (2018). In the case of superradiant phase transition, the effect of dissipation is previously explored Kirton and Keeling (2018). It is noteworthy to study the impact of dissipation in the ENET, which we explore in this manuscript. The Hamiltonian approach is insufficient to capture such dissipative dynamics. Hence, we adapt the Liouvillian approach motivated by the Lindblad quantum master equation to incorporate the non-unitary dynamics Breuer et al. (2002).
In this prescription, the Liouvillian gap ($\Delta$) plays an important role in characterizing the asymptotic convergence to the steady state Haga et al. (2021); Zhou et al. (2022); Prosen and Pižorn (2008); Kessler et al. (2012); Cai and Barthel (2013); Kastoryano and Eisert (2013); Žnidarič (2015); Casteels et al. (2017); Minganti et al. (2018); Shibata and Katsura (2019a, b, 2020); Mori and Shirai (2020, 2023); Shirai and Mori (2023); Mori (2024); Landi et al. (2022); Yuan et al. (2021), and one of our main focuses here is to study its dependence on the system size in different phases of this model. Recently, it was shown that for the boundary-dissipated one-dimensional quasiperiodic and disordered systems Zhou et al. (2022), $\Delta$ scales as exponentially for localized phase and $L^{-3}$ (where $L$ is the system size) for extended phase. For other types of boundary-dissipated systems, numerical results have shown that $\Delta$ scales with $L^{-z}$, where $z\in[1,2]$ for chaotic systems and $z=3$ for integrable systems  Prosen and Pižorn (2008); Žnidarič (2015); Shibata and Katsura (2020); Mori and Shirai (2020). On the other hand, for a special case ( XXZ chain in the presence of bulk dissipation ), it was found that, $\Delta$ is independent of the system size ($L$) for the ferromagnetic case, whereas it decays as $L^{-0.8}$ in the antiferromagnetic case Žnidarič (2015). Here, we study the behavior of $\Delta$ as a function of $L$ for dissipative ADM to characterize different phases. Moreover, we also investigate the properties of eigenvectors of the Liouvillian and compare their consistency with the corresponding eigenvalues.
@@
Another important aspect of a many-body quantum system is exploring the dynamics under an external drive Eckardt (2017). In such cases, it is expected that the system will reach the infinite temperature state (e.g., thermal state) Abanin et al. (2019); Nandkishore and Huse (2015). An important distinction can be observed for the periodic drive, where instead of thermalization, a prethermal quasi-steady state emerges in the intermediate time scale Abanin et al. (2015); Beatrez et al. (2021); Saha and Bhattacharyya (2023); Das et al. (2023b). Using the well-celebrated Floquet theory, an effective Hamiltonian can be constructed for the prethermal state in high-frequency limit and the relevant conserved quantities can be derived subsequently Fleckenstein and Bukov (2021). More recently, research interest shifted towards exploring the quantum dynamics under quasi-periodic drive Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022); Das et al. (2023b); Tiwari et al. (2024, 2025), where it is expected that the lifetime of the prethermal phase is shortened due to the absence of temporal order. As such, the quasi-periodic drive destroys the integrability of the system and pushes it to the thermal state. A similar study was also reported in our previous work of ADM Das et al. (2023b). In this manuscript, we extend this analysis to the dissipative case, where we study the interplay between the quasi-periodic drive and dissipation and its effect on the stability of the prethermal state.
The structure and the key points of the manuscript are discussed below. A brief description of the dissipative ADM (i.e., a spin-boson system interacts with another bosonic environment) is given in section-II. The dynamical equation of the system is given in Eq. (3). We also discuss the criticality in the system and the effect of dissipation on the critical line in section- II.1. In section- III, we study the scaling of $\Delta$ for our system, as a function of the total Liouvillian space dimension ($N_{\text{L}}$) and the atom number ($N$), separately for the nonergodic and the ergodic phase. Furthermore, we study the eigenvector property with the aid of the average participation ratio for left, right, and both left-right eigenvectors separately in the same section.
The effect of an external drive on dissipative ADM is provided in section IV, where we consider a quasiperiodic drive modeled by the Thue-Morse sequence Thue (1906); Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022) and study the quantum dynamics with the aid of average boson number and the mutual information between spins and bosons in the system. We mostly focus on the changes in the behavior of our model in comparison to its corresponding closed version, which has not been explored earlier. Finally, we conclude in section-V and discuss the broader perspective of our results.
Figure 1: Schematic picture for the dissipative Dicke model, where the system is defined by the interaction between an ensemble of $N$ two level atoms with a single mode ($\omega$) leaky bosonic cavity and the whole system is connected with another bosonic environment. The gap between the two levels of each atom is $\omega_{0}$ and the bosonic dissipation strength is $\kappa$.
@@ II Description of the system
where $g_{1}$ and $g_{2}$ are the rotating and counter-rotating spin-boson coupling strengths as we are interested in the anisotropic Dicke model Das et al. (2023a). Here $a$ ($a^{\dagger}$) is the bosonic annihilation (creation) operator, and $J_{\alpha}=\sum_{i=1}^{N}\frac{1}{2}\sigma_{\alpha}^{i}$ ($\alpha=$ $x$, $y$ $z$), total angular momentum of the ensemble of $N$ spins. $\omega$ is the bosonic mode frequency and $\omega_{0}$ is the level splitting of the atoms. The first two terms represent the free Hamiltonian of the bosonic cavity and atoms respectively. The basis of the Hamiltonian is: $|n,m\rangle$, where $n=0,1,2...,n_{\text{max}}$: boson number and for symmetric spin subspace considering even $J$, $m=-J,-J+1,...,0,...,J-1,J$. Thus the Hilbert space dimension of the closed system is: $N_{\text{D}}=(N+1)(n_{\text{max}}+1)$. $g_{1}=g_{2}=g$ recovers the original isotropic Dicke model.
We assume that the decay rate of the leaky cavity is $\kappa$ and all the other sources of dissipation and dephasing are excluded in our model. The coupling between the cavity and the environment is weak and it can be modeled using Jaynes-Cummings Hamiltonian Breuer et al. (2002). The presence of the environment leads to dissipative dynamics which can be explained by Lindblad master equation in the Markovian regime Breuer et al. (2002). In Fig. 1 we show a schematic diagram of the dissipative isotropic Dicke model.
The dynamical equation of atom and leaky cavity is given by:
@@
where $\rho$ is total density matrix of the total system and $\mathcal{L}$ is the Liouvillian super-operator. The explicit form of $\mathcal{L}\rho$ is:
@@
The basis of the Liouvillian is $|n_{l},m_{l}\rangle\otimes|n_{r},m_{r}\rangle$ and hence the Liouvillian space dimension is $N_{\text{L}}=N_{\text{D}}\times N_{\text{D}}$. Using parity symmetry, the dimension of the Liouvillian space reduces to $\frac{N_{\text{ L}}}{2}$ Prasad et al. (2022).
II.1 Existence of the critical points
Criticality in the closed Dicke model is well known Das and Sharma (2022) and in the presence of anisotropy, how this critical point is generalized to a critical line is also explored previously Das et al. (2023a). In this manuscript, one of our main goals is to understand the effect of dissipation on the critical line. For the bosonic dissipation, the relation between the critical coupling parameters for the second-order phase transition is given by Kirton et al. (2019):
@@ II.1 Existence of the critical points
which separates the non-ergodic and ergodic phases. In case of isotropic Dicke model ($g_{1}=g_{2}=g$), the equation will reduce to: $g^{*}=\frac{1}{2}\sqrt{\omega\omega_{0}}\sqrt{1+\kappa^{2}/\omega^{2}}$. For the numerical simplicity, we choose $\omega=\omega_{0}=1$. To include the dissipation we choose $\kappa=1$ and hence the critical point for the quantum phase transition from integrable to chaotic phase is: $g^{*}=\frac{1}{\sqrt{2}}\approx 0.7071$.
If we gradually increase the value of $\kappa$, then the regime of the non-ergodic phase will also be extended which means the non-ergodic phase is more favorable under the bosonic dissipation. In the next section, we quantify the above result by studying several properties of the Liouvillian.
III Results
III.1 Finite size scaling of the Liouvillian gap
Similar to the closed model, dissipative ADM also shows a transition from non-ergodic to ergodic phases (see Fig. 9 in the Appendix). In this section, we study the eigenvalue property of the Liouvillian $\mathcal{L}$, the scaling of the Liouvillian gap (as a function of the total Liouvillian space dimension) in different regions of the phase diagram covering the ergodic and non-ergodic phases.
Figure 2: The Liouvillian gap as a function of (a) the atom number $N$ and (b) the total dimension $N_{\text{L}}$ of the Liouvillian space for $g_{1}=1.25$, and $g_{2}=0.1-1.5$ such that we cover both the nonergodic and the ergodic phase of the system. For each pair of $(g_{1},\ g_{2})$ the range of the atom number is $N=4-10$. In the nonergodic phase $\Delta$ is independent of the system size whereas, in the ergodic phase: (a) $\Delta\propto N^{z}$. For, $g_{2}=1.0:\ \ln\Delta=-0.42[\ln N]-0.0435$, $g_{2}=1.1:\ \ln\Delta=-0.578[\ln N]+0.124$, $g_{2}=1.2:\ \ln\Delta=-0.75[\ln N]+0.28$, $g_{2}=1.25:\ \ln\Delta=-0.84[\ln N]+0.362$, $g_{2}=1.3:\ \ln\Delta=-0.897[\ln N]+0.38$, $g_{2}=1.4:\ \ln\Delta=-0.9945[\ln N]+0.384$, and $g_{2}=1.5:\ \ln\Delta=-1.053[\ln N]+0.34$; (b) $\Delta\propto N_{L}^{z}$ (in the ergodic phase). For $g_{2}=1.0:\ \ln\Delta=-0.244[\ln N_{L}]+1.76$, $g_{2}=1.1:\ \ln\Delta=-0.336[\ln N_{L}]+2.609$, $g_{2}=1.2:\ \ln\Delta=-0.445[\ln N_{L}]+3.61$, $g_{2}=1.25:\ \ln\Delta=-0.514[\ln N_{L}]+4.27$, $g_{2}=1.3:\ \ln\Delta=-0.55[\ln N_{L}]+4.56$, $g_{2}=1.4:\ \ln\Delta=-0.61[\ln N_{L}]+5.023$, and $g_{2}=1.5:\ \ln\Delta=-0.65[\in N_{L}]+5.26$. For all the plots $\omega=\omega_{0}=1$, $\kappa=1$. The bosonic cut-off: $n_{\text{max}}=26$.
Figure 3: Average of participation ratio ($P_{\text{avg}}$), using (a) the left (b) right, and (c) biorthonormal basis considering both left and right eigenvectors of the Liouvillian for dissipative ADM as a function of the rotating and counter-rotating coupling parameters $g_{1}$ and $g_{2}$. The atom number: $N=10$, the bosonic cut-off: $n_{\text{max}}=26$. For all the figures $\omega_{0}=\omega=1,\ \kappa=1$.
Figure 4: Average of participation ratio ($P_{\text{avg}}$), using both the left and right eigenvectors of the Liouvillian for dissipative DM as a function of the coupling parameter $g$. The atom number: $N=10$, the bosonic cut-off: $n_{\text{max}}=26$. $\omega=\omega_{0}=1$, $\kappa=1$.
The Liouvillian gap, $\Delta$ is defined by the real distance of the nearest neighbor eigenvalue of $\mathcal{L}$ from its zero eigenvalue (eigenvalue corresponds to the steady state):

### v2

# A detailed study on various phases in dissipative anisotropic Dicke model

@@ Abstract
We present a comprehensive study of different phases in the Dicke model incorporating both anisotropy and dissipation. We begin with a concise review of the quantum phase transition in this setting, highlighting how these two parameters shift the critical point. We then perform a detailed investigation of the transition from ergodic to nonergodic phases by analyzing the eigenvalue and eigenvector properties of the Liouvillian with the aid of scaling of the Liouvillian gap and the average participation ratio. Our results show that the eigenvector properties of the Liouvillian are consistent with its spectral characteristics, leading to a phase diagram that has similarities with the closed counterpart. Furthermore, we demonstrate that the Liouvillian gap exhibits distinct scaling behaviors in these two phases. Finally, we extend our study to the driven case by applying a Thue–Morse quasiperiodic drive. In this case, we find that bosonic dissipation plays a crucial role in stabilizing the prethermal plateau, offering an effective mechanism to halt the heating effect arising from the quasi-periodic drive.
@@ I Introduction
Exploring many-body quantum dynamics through atom-photon interaction is the heart of the atomic, condensed matter, and ultra-cold quantum physics Cohen-Tannoudji et al. (1998); Léonard et al. (2017); Gutzler et al. (2021); Chang et al. (2018). Experiments using cavity QED setup provide an ideal platform to investigate the underlying microscopic properties, e.g., the collective effects, various phase transitions, Bose-Einstein condensation, etc Walther et al. (2006); Reiserer and Rempe (2015); Vaidya et al. (2018); Dimer et al. (2007a). From a theoretical perspective, the Dicke model (DM) is an iconic model for understanding the above aspects of such systems Garraway (2011); Kirton et al. (2019); Villaseñor et al. (2024).
The study of collective light–matter interactions traces back to the pioneering work of Dicke, which describes the interaction between $N$ ensemble of atoms with a single-mode bosonic field via a dipolar coupling strength  Dicke (1954); Lambert et al. (2004); Emary and Brandes (2003a, b). In the thermodynamic limit ($N\to\infty$), the model shows a quantum phase transition (QPT) from normal to superradiant phase Dicke (1954); Chávez-Carlos et al. (2016); Kirton and Keeling (2018); Kirton et al. (2019); Das and Sharma (2022). Along with it, such a model also shows the excited state phase transition (ESQPT) and thermal phase transition (TPT) under different circumstances. First experimental proposal of observing the Dicke quantum phase transition using multilevel atoms and cavity-mediated Raman transitions were proposed by Dimer et al. Dimer et al. (2007b). Around the same time, Baumann et al. observed the Dicke phase transition in an open system consisting of a Bose-Einstein condensate coupled to an optical cavity, highlighting the emergence of a self-organized supersolid phase Baumann et al. (2010). Such studies later extended to observe Higgs and Goldstone modes in a supersolid quantum gas Léonard et al. (2017).
@@
A dissipative realization of the generalized Dicke model was also introduced to observe the effects from the counter-rotating terms, and the effect of external drive was further explored Gutiérrez-Jáuregui and Carmichael (2018a). The generalized version of the Dicke model, as mentioned above, namely the anisotropic Dicke model (ADM) Buijsman et al. (2017); Kloc et al. (2017); Gutiérrez-Jáuregui and Carmichael (2018b); Shapiro et al. (2020); Das et al. (2023a), where the coupling strengths correspond to the rotating and counter-rotating terms are considered to be different. Due to the asymmetry in the coupling strengths, ADM shows novel features in addition to the well-known existing properties, i.e., ergodic to non-ergodic transition (ENET) Buijsman et al. (2017); Hu and Wan (2021). There exists a one-to-one correspondence between the non-ergodic phase to the integrability of the system. Hence, a detailed study on the non-ergodic phase will be useful in understanding the following emerging phenomena, e.g., many-body localization, quantum scars, prethermalization, etc Polkovnikov et al. (2011); Mori et al. (2018); Serbyn et al. (2021). On the other hand, the ergodic phase connects to the non-integrable phase, a path to study the quantum chaos D’Alessio et al. (2016); Borgonovi et al. (2016). A natural question, therefore, concerns how the aforementioned phases behave in the presence of dissipation, and in particular whether dissipation destroys their non-ergodic character.
In most cases, a leaky cavity is considered to introduce the dissipative effects on the atom-photon interaction Bhaseen et al. (2012); Kirton et al. (2019); Kirton and Keeling (2017, 2018). In the case of QPT, the effect of dissipation is previously explored Kirton and Keeling (2018). It is noteworthy to study the impact of dissipation in the ENET, which we explore in this manuscript. The Hamiltonian approach is insufficient to capture such dissipative dynamics. Hence, we adapt the Liouvillian approach motivated by the Lindblad quantum master equation to incorporate the non-unitary dynamics Breuer et al. (2002). In this prescription, the Liouvillian gap ($\Delta$) plays an important role in characterizing the asymptotic convergence to the steady state Haga et al. (2021); Zhou et al. (2022); Prosen and Pižorn (2008); Kessler et al. (2012); Cai and Barthel (2013); Kastoryano and Eisert (2013); Žnidarič (2015); Casteels et al. (2017); Minganti et al. (2018); Shibata and Katsura (2019a, b, 2020); Mori and Shirai (2020, 2023); Shirai and Mori (2023); Mori (2024); Landi et al. (2022); Yuan et al. (2021). The dependence of such a Liouvillian gap on the system size for various systems has already been studied, where the nature of the dissipation and the corresponding phases determine such a scaling Zhou et al. (2022); Prosen and Pižorn (2008); Žnidarič (2015); Shibata and Katsura (2020); Mori and Shirai (2020). Based on these results, one can conjecture that the length dependence of the Liouvillian gap is also sensitive to the specific phases of many-body dissipative systems. Motivated by this perspective, we aim to investigate the system-size dependence of the Liouvillian gap for ENET in the ADM. Similarly, the level statistics of the eigenvalues of the Liouvillian provide useful information regarding the phases. For example, in the case of ADM, the system follows the $2D$ Poisson statistics in the non-ergodic phase, whereas in the ergodic phase the system behaves as the Ginibre ensemble Ginibre (1965); Grobe et al. (1988); Markum et al. (1999); Haake (1991); Hamazaki et al. (2020). For completeness of the above analysis, the distribution of the eigenvector is further required. We note that for non-Hermitian systems, the correspondence between the eigenvalue-eigenvector property is quite different from its Hermitian counterpart. Motivated by this fact, we performed the study of the average participation ratio of dissipative ADM to provide further explanation for the emergence of ENET.
@@
Another important aspect of a many-body quantum system is exploring the dynamics under an external drive Eckardt (2017). In such cases, it is expected that the system will reach the infinite temperature state (e.g., thermal state) Abanin et al. (2019); Nandkishore and Huse (2015). An essential distinction can be observed for the periodic drive, where instead of thermalization, a prethermal quasi-steady state emerges in the intermediate time scale Abanin et al. (2015); Beatrez et al. (2021); Saha and Bhattacharyya (2023); Das et al. (2023b). Using the Floquet theory, an effective Hamiltonian can be constructed for the prethermal state in the high-frequency limit, and the relevant conserved quantities can be derived subsequently Fleckenstein and Bukov (2021). More recently, research interest has shifted towards exploring quantum dynamics under quasi-periodic drive Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022); Das et al. (2023b); Tiwari et al. (2024, 2025), where it is expected that the lifetime of the prethermal phase is shortened due to the absence of temporal order. As such, the quasi-periodic drive further pushes it to the thermal state. A similar study was also reported in our previous work of ADM Das et al. (2023b). A further question concerns what happens in the presence of dissipation. We want to specify the reason why such a study is important. In the case of dissipative ADM, we consider the spontaneous emission of the cavity modes. Therefore, the cavity and the atoms try to reach their respective ground state in the presence of strong dissipation. The ground state has the lowest entanglement entropy (EE). At the same point, the quasi-periodic drive increases the EE due to heating. We expect a competition between the strength of the quasi-periodic drive and the dissipation rate in this system. We mostly focus on whether the strong dissipation stabilizes the prethermal phase or the quasi-periodic drive dominates the heating process. We also show that such a mechanism is different for the ergodic and non-ergodic phases.
As stated above, the primary motivation of this manuscript is to investigate the impact of dissipation on the emergent phases of the ADM in both undriven and driven scenarios. To this end, we outline the structure and main results of the paper as follows. In Sec . II, we introduce the dissipative ADM, where a spin–boson system is coupled to an additional bosonic environment. We begin with the undriven case in Sec. III. The existence of QPT is analyzed in Sec. III.1, followed by a detailed study of the ENET in Sec. III.2. In particular, we examine the Liouvillian gap in Sec. III.2.1 and investigate the properties of the corresponding eigenvectors in Sec. III.2.2. We then turn to the driven case in Sec. IV, where the quasi-periodic driving protocol is defined in Sec. IV.1, and the role of dissipation in the presence of drive is discussed in Sec. IV.2. Finally, we summarize our findings and present a broader perspective on their implications in Sec. V.
@@ II Description of the system
where $g_{1}$ and $g_{2}$ are the corresponding rotating and counter-rotating spin-boson coupling strength. In general, $g_{1}\neq g_{2}$ indicates the anisotropy in the Dicke model Das et al. (2023a), and $g_{1}=g_{2}=g$ recovers the original isotropic Dicke model. Here $a$ ($a^{\dagger}$) is the bosonic annihilation (creation) operator, and $J_{\alpha}=\sum_{i=1}^{N}\frac{1}{2}\sigma_{\alpha}^{i}$ ($\alpha=$ $x$, $y$ $z$), total angular momentum of the ensemble of $N$ atoms. $\omega$ is the bosonic mode frequency and $\omega_{0}$ is the level splitting of the atoms. The first two terms represent the free Hamiltonian of the bosonic cavity and atoms, respectively. The basis of the Hamiltonian is: $|n,m\rangle$, where $n=0,1,2...,n_{\text{max}}$: boson number and for the symmetric spin subspace considering even $J$, m=J,J+1,,0,,J1,Jm=-J,-J+1,...,0,...,J-1,J. Thus, the Hilbert space dimension of the closed system is: $N_{\text{D}}=(N+1)(n_{\text{max}}+1)$.
We assume that the spontaneous emission rate of the leaky cavity is $\kappa$ and all other sources of dissipation and dephasing are excluded in our model. The dynamics can be well explained by the Lindblad master equation in the Markovian regime Breuer et al. (2002). The dynamical equation of atom+cavity is given by:
@@
where $\rho$ is density matrix of the system and $\mathcal{L}$ is the Liouvillian super-operator. The explicit form of $\mathcal{L}\rho$ is:
@@
The basis of the Liouvillian is $|n_{l},m_{l}\rangle\otimes|n_{r},m_{r}\rangle$ and the corresponding Liouvillian space dimension is $N_{\text{L}}=N_{\text{D}}\times N_{\text{D}}$. Using parity symmetry, the dimension of the Liouvillian space is reduced to $\frac{N_{\text{ L}}}{2}$ Prasad et al. (2022). In such systems, we are primarily interested in the following phases, as summarized below.
  1. 1.
    Absence of external drive,
    1. (a)
      Normal and Superradiant phases
    2. (b)
      Ergodic and non-Ergodic phases
  2. 2.
    Presence of external drive,
    1. (a)
      Prethermal phase and further heating
In the remainder of the manuscript, we will briefly discuss the existence of such phases in ADM.
III Analysis for the undriven case
We begin with the undriven case, where we aim to study the effects of dissipation on QPT and ENET.
III.1 Emergence of quantum phase transition (QPT)
As QPT is well defined and has been discussed thoroughly in previous works, we provide a concise review of these phenomena. Let us consider an isotropic Dicke model, $g_{1}=g_{2}$. The Hamiltonian of the system is invariant under the following transformation, $a\to-a$, $\sigma_{x}\to-\sigma_{x}$ (i.e., $\mathbb{Z}_{2}$ symmetry). QPT is defined in this case as the breaking of $\mathbb{Z}_{2}$ symmetry. As such, the normal phase is the $\mathbb{Z}_{2}$ protected phase, with $\langle a\rangle=0$, and in the case of a symmetry-broken superradiant phase, $\langle a\rangle\neq 0$ Kirton et al. (2019). Following the retarded Green’s function and the Keldysh path integral formalism, the critical coupling strength for the second-order QPT is defined as $g_{c}=\frac{1}{2}\sqrt{\omega\omega_{0}}\sqrt{1+\kappa^{2}/\omega^{2}}$. Following the same procedure, in case of ADM, instead of a line, we get a region of $g_{1}$ and $g_{2}$, which satisfy the following relation,
@@ III.1 Emergence of quantum phase transition (QPT)
Several experimental platforms have been developed to observe QPT in the Dicke model Dimer et al. (2007b); Baumann et al. (2010); Léonard et al. (2017).
III.2 Emergence of ergodic to non-ergodic phase transition (ENET)
In this section, we provide a concise review of the ENET in closed ADM. In case of ADM, the system is integrable for $g_{1}=0$ or $g_{2}=0$. As a consequence, the system shows ENET in the vicinity of the integrable regime. Previous work shows the emergence of ENET in closed ADM with the aid of level spacing ratio and also shows that such a phase transition is not the precursor of QPT Buijsman et al. (2017). We will further extend this approach to the dissipative case and study the properties of the Liouvillian of such systems for different values of $g_{1}$ and $g_{2}$.
III.2.1 Finite size scaling of the Liouvillian gap
An analysis using the eigenvalues of the Liouvillian ($\mathcal{L}$) is required to understand ENET in ADM. The real part of the eigenvalues of $\mathcal{L}$ is always less than or equal to zero. The real zero eigenvalue shows the notion of the steady state. On the other hand, the negative real parts are the decaying modes. If the complex part is non-zero, then such eigenmodes show oscillatory behavior Saha and Bhattacharyya (2024); Kessler et al. (2012). Whether such oscillation is long-lived or not is decided by the corresponding real part of the eigenvalue. The Liouvillian gap, $\Delta$, is defined by the real distance of the nearest neighbor eigenvalue of $\mathcal{L}$ from its zero eigenvalue.